# Semantic Chunking and the Entropy of Natural Language

### v1

@@ I Introduction
Reading and understanding a natural text involves gradual comprehension of its meaning at different levels of abstraction, from single words to specific details, and finally to its overall gist. This hierarchical structure of texts generates strong statistical redundancies at all levels: words in a sentence have to satisfy the grammar, sentences have to relate to each other in a particular way, and the end of the story has to correspond to how the story begins. A reasonable, if necessarily simplified, way to describe the hierarchical structure of texts is by “semantic trees,” where each node corresponds to a particular “keypoint” that summarizes a certain part of the text [1, 2, 3, 4]. Every text will be characterized by its unique tree as its meaning develops, with different people forming different trees for the same text depending on how they comprehend it. In our recent publication [4], we argued that when considering multiple texts of a particular length, the corresponding ensemble of trees can be approximated by a random recursive partition of texts into shorter and shorter chunks. Moreover, we assumed that the splitting parameter of this ensemble, namely the maximal number of children for each node, is limited by working memory capacity.
@@
The entropy of natural language is inherently related to its semantic structure. The redundancy of natural language implies that one can often anticipate an upcoming word from its context. Crucially, this anticipation is not a single-step guess but operates through a hierarchy of abstract inferences: a reader first infers the global topic and communicative intent, then the local discourse function of the current section and paragraph, and only subsequently progressively refines expectations down to the immediate grammatical structure that determines the next word [5, 6, 7, 8]. This nested cascade of inferences is mirrored in the structure of the text itself in the form of semantic trees. These semantic chunks encode information about possible continuations before any word- or token-level realization is specified. Accordingly, the information captured by the semantic structure constitutes a component of the total entropy and can be viewed as a lower bound on the full entropy rate of language.
@@
On the other hand, the redundancy of language can be quantified by its entropy rate, relative to the maximal possible rate that depends on the size of the alphabet (i.e., the number of characters or words). The first estimate of the entropy rate of printed English was obtained by Shannon using the now-famous “guessing” game procedure, where a subject was trying to guess the next character after reading passages of text of various lengths [9, 10]. The resulting estimate of about 1 bit per character was then confirmed in other languages as well [11, 12]. Current large language models (LLMs) can now be used to estimate entropy rates of language without using human subjects and result in similar estimates [13, 14, 15, 16]. Until now, however, there is no first-principles understanding of what accounts for the entropy rate of English and other languages. The goals of this study are two-fold. First, we apply LLMs to segment multiple texts from several corpora, ranging from children’s books to modern poetry. We compare the obtained set of trees to a random ensemble proposed in [4] and find the splitting parameter that yields the best match in terms of chunk-size distribution. We then develop a theoretical framework for calculating the entropy of the random tree ensemble and compare the results with empirically obtained entropy rates across different corpora. Finally, we propose that the entropy rate of a particular text correlates with its comprehension difficulty and suggest future experiments to test this hypothesis.
@@ IV.0.2 Large $N$ limit
(a) Entropy rate as a function of $K$. Theoretical values correspond to $h_{K}$ in Eq. (11). Empirical values are obtained by selecting the optimal $K$ for each corpus (Table 1) and comparing $h_{K}$ to the corpus-level LLM estimate $h_{\mathrm{LLM}}$. Shaded bands indicate three approximate entropy-rate regimes that differentiate genres. (b) Entropy-rate estimates from individual random-tree realizations (simulated using Eq. (2)) concentrate around the predicted value as $N$ increases, consistent with the emergence of typical trees. (c) Per-token information for 100 RedditStories computed in two ways: from LLM perplexity and from the likelihood of empirical semantic trees obtained via chunking. As $N$ increases, both estimates fluctuate around the predicted entropy rate. (d)–(f) Cumulative LLM surprisal $\sum_{i=1}^{N}\log P(t_{i}|t_{<i})$ for 100 texts from each corpus (color indicates text length $N$). The blue dash-dotted fit gives the LLM-based entropy-rate estimate $h_{\mathrm{LLM}}$. The red dashed line shows the theoretical prediction $h_{K}$ using the optimal $K$ for each corpus; its intercept is chosen for visual comparison (matched to the blue fit).
@@ V From Semantic Trees to Entropy
$-\frac{1}{N}\log P(T)\xrightarrow{N\to\infty}h_{K}$ (12)
i.e. for large $N$ most of the trees generated according to the above procedure belong to a set of “typical trees” with per-leaf log-likelihooeds concentrated around entropy rate; total probability of this set approaches $1$ in this limit. This concentration is illustrated in Fig. 3(b): simulations at increasing $N$ show that realization-to-realization fluctuations in the estimated entropy rate vanish, converging to the theoretical value $h_{K}\approx 2.5$ nats/leaf for $K=4$.
@@ V.0.1 Selecting $K$ for each corpus
Furthermore, the inferred $K^{\star}$ aligns with intuitive notions of textual complexity (Fig. 3(a)). Simpler corpora such as children’s stories (red band; TinyStories and Fairytales) exhibit lower entropy rates and smaller $K^{\star}$ than narrative and expository corpora (green band; RedditStories and arXivAbstracts), which in turn fall below more stylistically atypical corpora such as poetry (blue band; ModernPoetry). Thus, both the entropy rate and the optimal branching factor suggest systematic differences across genres.
@@ VI Scaling in $L$ and universality of trees
the level-dependent scaling functions $f_{L}$ should collapse onto a universal standard normal distribution, $\mathcal{N}(0,1)$, independent of $L$ (and, after standardization, independent of $K$ as well). Consistent with this prediction, the functions $f_{L}$ differ substantially across levels when plotted in the original normalized-size coordinate $s_{L}$ (Fig. 2(e)). However, when replotted as functions of $x_{L}$, the same curves collapse onto $\mathcal{N}(0,1)$ as $L$ increases (Fig. 2(f)). We observe the same behavior empirically. The level-dependent empirical distributions in Fig. 2(g) collapse onto the same universal curve when expressed in the standardized coordinate $x_{L}=(\ln s_{L}-\hat{\mu}_{L})/\hat{\sigma}_{L}$, where $\hat{\mu}_{L}$ and $\hat{\sigma}_{L}$ are the empirical mean and standard deviation of $\ln s_{L}$ at level $L$ (Fig. 2(h)).
@@ III.2.1 Exact entropy rate for $K=2$
This recurrence relation has the same form as the one governing the average number of comparisons in the ’quicksort’ array sorting algorithm, which can be solved exactly using a differencing trick, as shown by Knuth [57]. Cover and Thomas were the first to apply the same technique to calculate the entropy of random binary trees, given as an exercise in their seminal book on information theory (Problem 4.5 in [58]). While their exercise corresponds to random strong integer ordered partition with $K=2$, here we apply it to our weak ordered partition random tree model with $K=2$. Despite this difference in the underlying combinatorics, the form of the recurrence remains the same, and thus can be solved using the same differencing trick.

### v2

@@ I Introduction
Reading and understanding a natural text involves gradual comprehension of its meaning at different levels of abstraction, from single words to specific details, and finally to its overall gist. This hierarchical structure of texts generates strong statistical redundancies at all levels: words in a sentence have to satisfy the grammar, sentences have to relate to each other in a particular way, and the end of the story has to correspond to how the story begins. A reasonable, if necessarily simplified, way to describe the hierarchical structure of texts is by “semantic trees,” where each node corresponds to a particular “keypoint” that summarizes a certain part of the text [1, 2, 3, 4]. Every text will be characterized by its unique tree as its meaning develops, with different people forming different trees for the same text depending on how they comprehend it. In our recent publication [4], we argued that when considering multiple texts of a particular length, the corresponding ensemble of trees can be approximated by a random recursive partition of texts into increasingly smaller chunks. Moreover, we assumed that the splitting parameter of this ensemble, namely the maximal number of children for each parent chunk, is limited by the human working memory capacity.
@@
This hierarchical semantic organization is closely tied to the predictability of language. The redundancy of natural language implies that one can often anticipate an upcoming word from its context. Crucially, this anticipation is not a single-step guess but operates through a hierarchy of abstract inferences: a reader first infers the global topic and communicative intent, then the local discourse function of the current section and paragraph, and only subsequently progressively refines expectations down to the immediate grammatical structure that determines the next word [5, 6, 7, 8]. This nested cascade of inferences is mirrored in the structure of the text itself in the form of semantic trees. These semantic chunks encode information about possible continuations before any word- or token-level realization is specified. Accordingly, the information captured by the semantic structure should account for a principled component of a text’s uncertainty.
@@
Historically, this uncertainty has been quantified by the entropy rate of language, a measure of information content relative to the maximal possible rate determined by the size of the alphabet (i.e., the number of characters or words). The first estimate of the entropy rate of printed English was obtained by Shannon using the now-famous “guessing” game procedure, where a subject was trying to guess the next character after reading passages of various lengths [9, 10]. The resulting estimate of approximately 1 bit per character, which implies 80% redundancy relative to the 5 bits per character of random text, was subsequently confirmed in other languages [11, 12]. Modern large language models (LLMs) can now be used to estimate entropy rates of language without using human subjects and result in similar estimates [13, 14, 15, 16].
Until now, however, there is no first-principles understanding of what accounts for the observed entropy rate of English and other languages. In this work, we derive this entropy rate directly from the hierarchical semantic organization of language. We first demonstrate that the random tree ensemble theory [4] accurately predicts the statistics of semantic trees obtained from the recursive segmentation of texts. We then develop a theoretical framework to calculate the entropy of this ensemble, showing that our theoretical predictions closely match empirical entropy rates measured by LLMs across diverse corpora. Crucially, our theory relies on a single free parameter $K$ representing the maximum branching factor (number of chunks) at each level of the tree. Remarkably, setting $K=4$ recovers Shannon’s classic estimate of $\sim$ 1 bit per character as a special case. While regular texts (e.g., novels and arXiv paper abstracts) typically aligns with this rate, our theory reveals that the entropy rate is not a fixed constant but increases systematically with text complexity, ranging from children’s books ($K\approx 2$) to modern poetry ($K\approx 6$). Finally, we propose that the entropy rate serves as a quantifiable proxy for comprehension difficulty and outline future experiments to test this hypothesis.
@@ IV.0.2 Large $N$ limit
(a) Entropy rate as a function of $K$. Theoretical values correspond to $h_{K}$ in Eq. (11). Empirical values are obtained by selecting the optimal $K$ for each corpus (Table 1) and comparing $h_{K}$ to the corpus-level LLM estimate $h_{\mathrm{LLM}}$. Shaded bands indicate three approximate entropy-rate regimes that differentiate genres. Shannon’s original estimate of 1 bit per character corresponds to approximately 2.2-2.8 nats per token (assuming 3-4 characters per token), lies near $K=4$ on this curve. (b) Entropy-rate estimates from individual random-tree realizations (simulated using Eq. (2)) concentrate around the predicted value as $N$ increases, consistent with the emergence of typical trees. (c) Per-token information for 100 RedditStories computed in two ways: from LLM perplexity and from the likelihood of empirical semantic trees obtained via chunking. As $N$ increases, both estimates fluctuate around the predicted entropy rate. (d)–(f) Cumulative LLM surprisal $\sum_{i=1}^{N}\log P(t_{i}|t_{<i})$ for 100 texts from each corpus (color indicates text length $N$). The blue dash-dotted fit gives the LLM-based entropy-rate estimate $h_{\mathrm{LLM}}$. The red dashed line shows the theoretical prediction $h_{K}$ using the optimal $K$ for each corpus; its intercept is chosen for visual comparison (matched to the blue fit).
@@ V From Semantic Trees to Entropy
$-\frac{1}{N}\log P(T)\xrightarrow{N\to\infty}h_{K},$ (12)
i.e. for large $N$ most of the trees generated according to the above procedure belong to a set of “typical trees” with per-leaf log-likelihoods concentrated around entropy rate $h_{K}$; the total probability of this set approaches $1$ in this limit. This concentration is illustrated in Fig. 3(b): simulations at increasing $N$ show that realization-to-realization fluctuations in the estimated entropy rate vanish, converging to the theoretical value $h_{K}\approx 2.5$ nats/leaf for $K=4$.
@@ V.0.1 Selecting $K$ for each corpus
Notably, the original Shannon estimate of 1 bit per character, corresponding to approximately 2.2–2.8 nats per token (assuming 3–4 characters per token), aligns with $K=4$ in our theoretical prediction (Fig. 3(a)). Furthermore, the inferred $K^{\star}$ aligns with intuitive notions of textual complexity (Fig. 3(a)). Simpler corpora such as children’s stories (red band; TinyStories and Fairytales) exhibit lower entropy rates and smaller $K^{\star}$ than narrative and expository corpora (green band; RedditStories and arXivAbstracts), which in turn fall below more stylistically atypical corpora such as poetry (blue band; ModernPoetry). Thus, both the entropy rate and the optimal branching factor suggest systematic differences across genres.
@@ VI Scaling in $L$ and universality of trees
the level-dependent scaling functions $f_{L}$ should collapse onto a universal standard normal distribution, $\mathcal{N}(0,1)$, independent of $L$ (and, after standardization, independent of $K$ as well). Consistent with this prediction, the functions $f_{L}$ differ substantially across levels when plotted in the original normalized-size coordinate $s_{L}$ (Fig. 4(a)). However, when replotted as functions of $x_{L}$, the same curves collapse onto $\mathcal{N}(0,1)$ as $L$ increases (Fig. 4(b)). We observe the same behavior empirically. The level-dependent empirical distributions in Fig. 4(c) collapse onto the same universal curve when expressed in the standardized coordinate $x_{L}=(\ln s_{L}-\hat{\mu}_{L})/\hat{\sigma}_{L}$, where $\hat{\mu}_{L}$ and $\hat{\sigma}_{L}$ are the empirical mean and standard deviation of $\ln s_{L}$ at level $L$ (Fig. 4(d)).
@@ III.2.1 Exact entropy rate for $K=2$
This recurrence relation has the same form as the one governing the average number of comparisons in the “quicksort” array sorting algorithm, which can be solved exactly using a differencing trick, as shown by Knuth [57]. Cover and Thomas were the first to apply the same technique to calculate the entropy of random binary trees, given as an exercise in their seminal book on information theory (Problem 4.5 in [58]). While their exercise corresponds to random strong integer ordered partition with $K=2$, here we apply it to our weak ordered partition random tree model with $K=2$. Despite this difference in the underlying combinatorics, the form of the recurrence remains the same, and thus can be solved using the same differencing trick.